# Sample size and power calculation for propensity score analysis of observational studies

### v2

@@ A.4. Proof of Proposition 3
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (63)
$=\int u^{2}\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (64)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (65)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (66)
@@
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=a_{0}^{2}\sigma_{e}^{2}+\sigma_{0}^{2}+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (67)
Adding (57) and (67), we have
$V=(a_{1}^{2}+a_{0}^{2})\sigma_{e}^{2}+(\sigma_{1}^{2}+\sigma_{0}^{2})+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right)+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (68)
A.7. Proof of the asymptotic variance of the general Hájek WATE estimator.
@@
$\phi=\begin{pmatrix}\phi_{1}\\ \phi_{0}\\ \psi_{\alpha}\end{pmatrix}\quad\text{with}\quad\begin{aligned} \phi_{1}&=Z_{i}\cdot w_{1}(X_{i};\alpha)\cdot\{Y_{i}(1)-\xi_{1}\},\\ \phi_{0}&=(1-Z_{i})\cdot w_{0}(X_{i};\alpha)\cdot\{Y_{i}(0)-\xi_{0}\},\\ \psi_{\alpha}&=e^{\prime}(X_{i};\alpha)\{Z_{i}-e(X_{i};\alpha)\}.\end{aligned}$ (69)
@@
$\sqrt{N}\begin{pmatrix}\hat{\xi}_{1}^{-}\xi_{1}^{*}\\ \hat{\xi}_{0}^{-}\xi_{0}^{*}\\ \hat{\alpha}-\alpha^{*}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,(A^{*})^{-1}B^{*}(A^{*})^{-\top}),$ (70)
@@
$\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\big{|}_{(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})}=\begin{pmatrix}\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}&0&\frac{\partial\phi_{1}^{*}}{\partial\alpha^{\top}}\\ 0&\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}&\frac{\partial\phi_{0}^{*}}{\partial\alpha^{\top}}\\ 0&0&\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha^{\top}},\end{pmatrix}$ (71)
@@
A.8. Proof of (24)
@@
$\xi_{z}=\frac{\mathbb{E}[h(X)Y(z)]}{\mathbb{E}[h(X)]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})\epsilon_{z}]}{\mathbb{E}[h(W_{e})]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\mu_{z}.$ (72)
@@
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (73)
$=\mathbb{E}\left[\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\mid W_{e}\right]\right]$ (74)
$=\mathbb{E}\left[\left\{a_{1}^{2}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}+\sigma_{1}^{2}\right\}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (75)
$=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (76)
@@
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (77)
Appendix B Examples
B.1. Normality of linear combination of covariates.
@@ Appendix B Examples
B.2. Examples of the Beta approximation of logit-normal distribution.
Figure 5 presents the density of the Beta distribution over a grid of $(\alpha,\beta)$ values within $\{0.5,1,2,3\}$, and the density of the corresponding approximated logit-normal distributions.
Figure 5: Examples of Beta approximation to logit-normal distribution
@@ B.3 Calculation of overlap coefficient $\phi$ with fitted propensity scores
$\phi=\int_{0}^{1}\sqrt{f_{0}(u)f_{1}(u)}\,\mathrm{d}u=\frac{\int_{0}^{1}\sqrt{u(1-u)}\Pr(e(X)=u)\,\mathrm{d}u}{\sqrt{r(1-r)}}.$ (79)
@@
$\phi=\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}F(u)}{\sqrt{r(1-r)}}\approx\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}\widehat{F}(u)}{\sqrt{r(1-r)}}=\frac{1}{n}\cdot\frac{\sum_{i=1}^{n}\sqrt{\widehat{e}_{i}(1-\widehat{e}_{i})}}{\sqrt{r(1-r)}}.$ (80)
Appendix C. Additional Simulation and RHC Study Results
C.1. Density plot of the estimated propensity scores in the RHC study.
@@ Appendix C. Additional Simulation and RHC Study Results
C.2. Summary quantities in Simulation 5.1
$\kappa$ 0 0.25 0.5 0.75 0.9 1

### v3

@@ A.4 Proof of Proposition 2
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.38)
$=\int u^{2}\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.39)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.40)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.41)
@@
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=a_{0}^{2}\sigma_{e}^{2}+\sigma_{0}^{2}+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.42)
Adding (A.32) and (A.42), we have
$\small V=(a_{1}^{2}+a_{0}^{2})\sigma_{e}^{2}+(\sigma_{1}^{2}+\sigma_{0}^{2})+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right)+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.43)
A.7 Proof that the calculated sample size increases with $\rho^{2}$.
We rewrite $\widetilde{V}=\widetilde{V}_{\mathrm{SH}}+\rho^{2}\widetilde{V}_{\mathrm{adj}}$, where $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$ and
$\widetilde{V}_{\mathrm{adj}}=\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}-2+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}-1\right]\exp\left(-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}-1\right]\exp\left(\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right).$ (A.44)
Recall that $W\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$. Then, $\sigma_{e\mid z}^{2}=\mathbb{V}[W\mid Z=z]\leq\mathbb{V}[W]=\sigma_{e}^{2}$. Hence,
$\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}\geq 1,\quad\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid z}^{2}}-1\geq\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}-1\geq 0.$ (A.45)
Therefore, $\widetilde{V}_{\mathrm{adj}}\geq 0$.
A.8 Proof of the asymptotic variance of the general Hájek WATE estimator.
@@
$\phi=\begin{pmatrix}\phi_{1}\\ \phi_{0}\\ \psi_{\alpha}\end{pmatrix}\quad\text{with}\quad\begin{aligned} \phi_{1}&=Z_{i}\cdot w_{1}(X_{i};\alpha)\cdot\{Y_{i}(1)-\xi_{1}\},\\ \phi_{0}&=(1-Z_{i})\cdot w_{0}(X_{i};\alpha)\cdot\{Y_{i}(0)-\xi_{0}\},\\ \psi_{\alpha}&=e^{\prime}(X_{i};\alpha)\{Z_{i}-e(X_{i};\alpha)\}.\end{aligned}$ (A.46)
@@
$\sqrt{N}\begin{pmatrix}\hat{\xi}_{1}^{-}\xi_{1}^{*}\\ \hat{\xi}_{0}^{-}\xi_{0}^{*}\\ \hat{\alpha}-\alpha^{*}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,(A^{*})^{-1}B^{*}(A^{*})^{-\top}),$ (A.47)
@@
$\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\big{|}_{(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})}=\begin{pmatrix}\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}&0&\frac{\partial\phi_{1}^{*}}{\partial\alpha^{\top}}\\ 0&\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}&\frac{\partial\phi_{0}^{*}}{\partial\alpha^{\top}}\\ 0&0&\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha^{\top}},\end{pmatrix}$ (A.48)
@@
When the propensity scores are assumed known, $e^{\prime}(X_{i};\alpha^{*})=0$, and hence $a_{13}=a_{23}=0$. Consequently, the variance reduces to $a_{11}^{-2}(b_{11}+b_{22})$. However, unlike ATE, there is no theoretical guarantee on whether assuming known propensity score leads to a larger or a smaller variance.
A.9 Proof of (C.5)
@@
$\xi_{z}=\frac{\mathbb{E}[h(X)Y(z)]}{\mathbb{E}[h(X)]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})\epsilon_{z}]}{\mathbb{E}[h(W_{e})]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\mu_{z}.$ (A.49)
@@
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.50)
$=\mathbb{E}\left[\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\mid W_{e}\right]\right]$ (A.51)
$=\mathbb{E}\left[\left\{a_{1}^{2}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}+\sigma_{1}^{2}\right\}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.52)
$=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.53)
@@
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.54)
B Examples
B.1 Normality of linear combination of covariates.
@@ B Examples
B.2 Examples of the Beta approximation of logit-normal distribution.
Figure 5 presents the density of the Beta distribution over a grid of $(a,b)$ values within $\{0.5,1,2,3\}$, and the density of the corresponding approximated logit-normal distributions.
Figure 5: Examples of the Beta approximation to logit-normal distribution
@@ B.3 Calculation of overlap coefficient $\phi$ with fitted propensity scores
$\phi=\int_{0}^{1}\sqrt{f_{0}(u)f_{1}(u)}\,\mathrm{d}u=\frac{\int_{0}^{1}\sqrt{u(1-u)}\Pr(e(X)=u)\,\mathrm{d}u}{\sqrt{r(1-r)}}.$ (B.2)
@@
$\phi=\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}F(u)}{\sqrt{r(1-r)}}\approx\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}\widehat{F}(u)}{\sqrt{r(1-r)}}=\frac{1}{n}\cdot\frac{\sum_{i=1}^{n}\sqrt{\widehat{e}_{i}(1-\widehat{e}_{i})}}{\sqrt{r(1-r)}}.$ (B.3)
C Extension to weighted average treatment effect estimands
$\tau^{h}=\frac{\operatorname{\mathbb{E}}[\tau(x)h(x)]}{\operatorname{\mathbb{E}}[h(x)]}.$ (C.1)
The Hájek estimator for the WATE is:
$\widehat{\tau}_{w}=\frac{\sum_{i}w_{1}(X_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(X_{i})Z_{i}}-\frac{\sum_{i}w_{0}(X_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(X_{i})(1-Z_{i})}.$ (C.2)
$V_{w}\equiv\operatorname{\mathbb{V}}(\widehat{\tau}_{w})=\frac{1}{\mathbb{E}[h(X_{i})^{2}]}\mathbb{E}\left[\left\{\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}+\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}\right\}h(X_{i})^{2}\right],$ (C.3)
$\int_{-\infty}^{\infty}h(u)^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u,$ (C.4)
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.5)
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.6)
Similarly as the case for ATE, the result does not depend on $E_{1}$ or $E_{0}$, and only depends on $(r,\phi,\rho^{2})$ if one assumes $S_{1}^{2}=S_{0}^{2}$ and $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$.
D Additional Simulation and RHC Study Results
D.1 Density plot of the estimated propensity scores in the RHC study.
@@ D Additional Simulation and RHC Study Results
D.2 Summary quantities in Simulation 4.1
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00
$\phi$ 1.00 0.98 0.93 0.87 0.83 0.81
$r$ 0.50 0.50 0.50 0.50 0.50 0.50
$E_{1}$1.73 -1.90 -2.03 -2.13 -2.17 -2.19
$E_{0}$2.74 -2.56 -2.43 -2.34 -2.30 -2.27
$S_{1}^{2}$ 20.02 20.30 20.41 20.46 20.47 20.48
$S_{0}^{2}$ 20.03 19.70 19.46 19.28 19.20 19.14
$\rho_{1}^{2}$ 0.02 0.04 0.04 0.04 0.04 0.04
$\rho_{0}^{2}$ 0.03 0.03 0.02 0.02 0.02 0.02
$S^{2}$ 20.02 20.00 19.93 19.87 19.83 19.81
$\rho^{2}$ 0.02 0.03 0.03 0.02 0.02 0.02
$R^{2}$ 0.20 0.20 0.19 0.19 0.19 0.19
Table 2: Summary quantities $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, pooled summary quantities $(S^{2},\rho^{2})$ and the regression $R^{2}$ under different values of $\kappa$ in Simulation 4.1.
D.3 Simulations with binary outcomes
We repeat the simulation design as in Section 4.1, and dichotomize the potential outcome $Y_{i}(z)$ into a binary variable $Y^{b}_{i}(z)=\mathbb{I}(Y_{i}(z)>-2)$. The observed outcome $Y^{b}_{i}=Z_{i}Y^{b}_{i}(1)+(1-Z_{i})Y^{b}_{i}(0)$. The true ATE on $Y^{b}$ is $\tau^{b}=\mathbb{E}[Y^{b}_{i}(1)-Y^{b}_{i}(0)]=0.089$.
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00